OpenStudy (anonymous):

Explain, in complete sentences, the effect of the difference of squares pattern on the multiplication of radicals

OpenStudy (anonymous):

This example might help: \((1+\sqrt 2)(1-\sqrt 2)=1+\sqrt 2 - \sqrt 2 + 2 = 1+ 2 = 3\). Note how the final simplification has no radicals in it.